A carpenter is building a rectangular room with a fixed perimeter of 324 ft. What dimentions would yield the maximum area? What is the maximum area? I understand the formula P=2l+2w will help solve this, but am unsure how to find the length and width. Any help you can give is greatly appreciated. Thanks.
Answer by stanbon(57347)

A carpenter is building a rectangular room with a fixed perimeter of 324 ft. What dimentions would yield the maximum area? What is the maximum area?
-------------------
2(L + W) = 324
L+W = 162
L = (162-W)
---------------
Area = LW
Area = (162-W)W
A = 162W-W^2
----
You have a quadratic equation with a = -1, b = 162, c = -A
Maximum Area occurs when W = -b/(2a) = -162/(2*-1) = 81 ft.
--
Solve for "L":
L + W = 162
L = 162-81 = 81 ft.
-----
Maximum Area = 81^2 = 6561 sq. ft.
=====================================
Cheers,
Stan H.
=====================================